# Unified Software Engineering agent as AI Software Engineer

### v1

@@ 6.2. RQ2: Dissection of Plausible Patches
For USEagent, we manually investigated a sample of 218 plausible solutions (i.e. passing the resolution criteria of each benchmark) to understand the degree of memorization, overfitting or other anomalies. We report a low number of 3 cases of memorization (1.3%), next to 23 cases of overfitting (10.5%). The percentage of overfitting solutions is lower than the previously reported overfitting rate of 31% on SWE-bench from AutoCodeRover (Ruan et al., 2024). We attribute the lower overfitting rate in part to our extensive ExecuteTests action in USEagent, which allows for more versatile test execution. As a result, we see an overfitting rate of 13.6% from USEagent on the SWE-verified dataset. Another reason for the lower overfitting rate is from USEbench: the SWT and REPOTEST datasets show low numbers of overfitting, which further reduces the overall rate. Overfitting patches usually bypass tests using if-else for early return values. Sometimes, this matches the intended handling of edge-cases, whether such patches are overfitting was determined by distance to ground truth. Furthermore, we identified a set of 33 anomalies that are plausible, and do not overfit, yet they are suspiciously different from the ground truth. One example anomaly includes the introduction of a new vector library for xarray (a vector library itself). Another example is an unorthodox class-inheritance check for scikit-learn - instead of utilizing pythons standard function isinstance, a predicate is applied to the objects obj.__classes__. These solutions are functionally correct but are unlikely to be acceptable by developers, so we have marked them as anomalies.
Summary: RQ2 - Analysis of Plausible Solutions We manually inspected 218 plausible solutions from USEagent and report 10.5% overfitting, 1.3% memorization. Overfitting was most common within SWE-verified, and least common in REPOTEST and SWT.
@@ 6.3. RQ3: Self-Configuration
(a) Distribution of SWE Steps - the most common path starts with reproduction followed by code retrieval, editing and review.
(b) Distribution of SWT Steps - most common trajectory begins with test-case retrieval, test-editing and execution.
@@
Figure 4 shows the different choices of the Meta-Agent when facing different benchmarks (we present two for brevity). The histograms display the actions taken first, second, etc. including the termination action Terminate. We observe some clear patterns in the different benchmarks that match our intuition. For program repair tasks (i.e. in SWE-Verified), we see in Figure 4(a) that the first step prominently consists of generating a reproduction test, followed by retrieving relevant code context. Over the course of agent execution, the EditCode and ExecuteTests become more prominent, introducing and verifying code changes to the software. For regression testing (in SWT), Figure 4(b) shows that by-large the first step chosen is TestRetrieval, immediately followed by a (test-)code edit. The majority of USEagent trajectories on SWT tasks then consists of alternating test-changes and test-executions, as shown in the histograms. Although we observe a pattern in the sequence of action invocations, some invocations seem noisy. For example, Reproduction action is to generate a system-level reproducer test, but it is still invoked in some of the SWT trajectories (the requirement from SWT is to generate unit-level tests). We manually investigated some cases and found that the reproducers were later transformed into unit-level tests. Nevertheless, the sequence of action invocation follows a certain pattern for each task type as shown in Figure 4, showing that USEagent and self-configure its workflow based on the task type. Across all benchmarks, USEagent first gathers information (using the retrieval actions or reproducers), converging towards alternating EditCode and ExecuteTests until either budgets are exhausted or Terminate is chosen. This behavior showcases the capability of emulating concepts like Test-Time Scaling (Brown et al., 2024) and command-discovery (Bouzenia and Pradel, 2024) through the Meta-Agent.
Summary: RQ3 - Self Configuration USEagent is capable of choosing different action-patterns for different benchmarks: Program repair (SWE) starts with Reproduction, regression testing (SWT) with TestRetrieval. Initial steps generate information before the actions converge towards alternating EditCode and ExecuteTests until termination.
@@ 6.4. RQ4: Open Challenges in Agentic Systems
Based on our previous investigations, we identify several major challenges in the current unified software engineering agent. Firstly, when the task requires writing a large amount of code (e.g. feature development tasks), the agent often generates solutions that are “mostly correct”. This is reflected on our evaluation on the REPOCOD dataset, where the task is to implement a complex function based on natural language requirement, and the correctness is defined as passing a large number of hidden unit tests. We observe a number of solutions generated by the agent implement the required feature almost correctly, but fail a few hidden tests due to missing edge cases. The edge-case behaviors are sometimes not specified clearly in the natural-language task description. We attribute this challenge in part to the inherent ambiguity of natural language. A future agent may first resolve the ambiguity in natural language by transforming the task requirements into more “formal” artifacts such as specifications and test cases, and then generate solutions based on the formal specification. Another approach is to design human-agent interaction schemes to clarify ambiguity when needed.
@@
Secondly, the current agent lacks the capability of “backtracking” when the agent execution does not yield meaningful results on a given execution path. For example, when there is a partial patch at a particular program location, the agent is prone to continuously make small edits to the partial patch and execute tests to verify them. However, the agent may not make good progress with this approach, since it may be impossible to craft a good solution at the location of the partial patch. This issue is amplified in the SWETRY dataset, in which the task description comes with a partial patch. To tackle this issue, a future agent should employ a backtracking mechanism to discard unpromising partial solutions, and start over at a previous step. It could be possible to employ the recent reasoning LLMs to examine the agent execution trace and decide whether to backtrack, since reasoning models have shown backtracking behaviors in their thinking process (Guo et al., 2025). Moreover, recent work (Antoniades et al., 2024) on employing search algorithm on agent execution trajectories can also help the agent to escape from unpromising paths.
@@
Lastly, we still observe patch overfitting in tasks such as program repair and regression testing. Overfitting is a known problem for automated program repair (Goues et al., 2019). Overfitting happens due to the generated patches passing given tests, but missing actual requirements. Thus, to avoid overfitting, if the tests are too specific, we need to generalize them. Agentic systems show promise for test generation (also reflected in this work), and the framework introduced by USEagent allows for additional agents that cover test-amplification or mutation testing, adversarial reasoning or test carving. The produced artifacts will feed back into test-execution, code generation, review, etc. and reduce overfitting as well as increasing efficacy and trustworthiness.
@@
Summary: RQ4 - Open Challenges Challenges consist of exploring edge-case handling and alternative solutions in feature development, backtracking and discarding poor partial results in repair, and repair-overfitting. We suggest search-approaches for the first, and agents for test-amplification for the latter.
@@ 7. Related Work
Recent Software Engineering Benchmarks fall into two broad categories: Isolated coding exercises and repository level tasks. Some coding exercises like CodeXGlue (Lu et al., 2021) come without an evaluation metric and form a training corpus, others like HumanEval (Chen et al., 2021a), MBPP (Austin et al., 2021), ClassEval (Du et al., 2023) or CoderEval (Yu et al., 2024a) provide a challenge accompanied by an evaluation harness (i.E. a test-suite). There are already ongoing efforts in providing a meta-benchmark for these tasks (Yu et al., 2024b)(Liu et al., 2024), which commonly target LLMs directly, less agentic systems. For USEBench, we select repository-level tasks to provide challenges of a uniform granularity. We decided against coding exercises as many models slowly saturate on them (Sallou et al., 2024).
Large Language Models and Agentic Systems
We consider a system agentic if the LLMs have a certain autonomy and are able to interact with the system through means of commands or other pre-defined tools. As such, Agentless (Xia et al., 2024) utilizes LLMs but is no agentic system due to its pre-defined workflow. Within agentic systems we see a divide into raw LLM-Agents that interface LLMs directly with a console and minimal tooling (e.g. ReAct (Yao et al., 2023), Gemini (Google, 2024) or SWE-Agent (Yang et al., 2024)) and approaches that introduce more domain knowledge as well as additional techniques. These additions range from specialised tools (like reproduction and fault localisation in AutoCodeRover (Zhang et al., 2024b)) to self-reflective capabilities (like the reviewer-agent in SpecRover (Ruan et al., 2024)). For this work, we have focused on AutoCodeRover (Zhang et al., 2024b) and OpenHands (Wang et al., 2024b), as these are open source projects with big resonance in the community. While the AIDER project (AIDER Team, 2024) has a well-suited scope (multiple SE tasks), its workflow relies on human prompting. We position AutoDev (Tufano et al., 2024) similar to AIDER, as the primary interface is a CLI for human interaction.
Agentic Systems
Passerine (Rondon et al., 2025) by Google proposes a program repair framework oriented at SWEagent (Yang et al., 2024) incorporating a similar structure to our meta-agent, providing a total of 5 pre-configured commands (excluding free commandline). Our efforts differ by scope as we target more than program repair, and employ a larger set of agents to multiple tasks. Similarly, Microsoft’s MASAI (Wadhwa et al., 2024), defines an action-space for their central agent similar to the meta-agent, limited to program repair. We differ from both works by providing more agents, structure the systems’ information and responses through a task-state, and apply the agentic system to more tasks than program repair. CodeR (Chen et al., 2024) has a meta-agent-component that pre-defines an execution plan, but does not adjust it dynamically.
@@ 8. Threats to validity
Within USEbench, we approximate testing capabilities exploiting test-coverage. Test-coverage does not imply a semantic coverage of the code, and can be achieved by various hacks. In lack of a better oracle we have opted for test-coverage as it still captures many of the attributes we want to evaluate: The patches must be in a correct format, location and fit into the projects test-suite. Furthermore, upon inspection many of the REPOTEST datapoints contain branches that handle edge-case behavior with errors, for which a syntactic coverage equals semantic testing.
Internal Validity
Data Leakage It is possible that the LLMs have been exposed to Code subject to evaluation or even datapoints from the source-benchmarks(Sallou et al., 2024). This is particularly likely as many publications (Yang et al., 2024; Ruan et al., 2024; Xia et al., 2024) utilize the APIs by common providers that use user-data for training. Combating this is beyond our capabilities, but we aim to provide an overview of memorization by manually inspecting a sample of positives.
@@ 9. Discussion
While we consider that costs will play a diminishing role for agentic systems as foundation models and infrastructure advances, we still want summarize the accumulated costs over the experiments. USEagent costs were at an average of 3.65 USD, with notable differences between sub-benchmarks. Herein SWE scored lowest with 2.20 USD average, and REPOCOD highest at a mean of 7.66 USD. The high costs at REPOCOD originate from the iteration of patch-generation and test-execution. Within actions, EditCode was the dominant cost-driver, due to the (large) contexts provided through relevant code and tests. We calculate the Pearson correlation coefficient for the elapsed time and the induced costs, and find a strong correlation of $r=0.89$. This shows a strong correlation between computational (e.g. test execution) and model (e.g. reasoning) efforts.
@@
We consider the approach of USEagent in combination with the meta-benchmark USEbench to be a promising avenue for any task that revolves around code as its primary artifact. Many tasks such as dependency management or project configuration can be incrementally tackled by extending USEbench and introducing more actions into USEagent. For USEagent to become truly an AI Software Engineer, it must address tasks such as requirements engineering, data visualization, deployment, code review or even a-b-testing. The efforts presented in this work are the second step in agentic development towards an AI software engineer, the first being individual agentic systems for single tasks (like AutoCodeRover). With the USEagent we pave the way for a unified approach to interact autonomously with code regardless of the task, allowing us to extend our scope beyond its current limits. Going beyond the USEagent will involve studying the AI-AI and AI-human collaboration. Thus we will need to study the cooperative intelligence resulting from multiple USEagents and multiple human developers. This might establish the dynamics of future development teams.

### v2

@@ 6.2. RQ2: Dissection of Plausible Patches
For USEagent, we manually investigated a sample of 218 plausible solutions (i.e. passing the resolution criteria of each benchmark) to understand the degree of memorization, overfitting or other anomalies. We report a low number of 3 cases of memorization (1.3%), next to 23 cases of overfitting (10.5%). The percentage of overfitting solutions is lower than the previously reported overfitting rate of 31% on SWE-bench from SpecRover (ruan2024specrover). We attribute the lower overfitting rate in part to our extensive ExecuteTests action in USEagent, which allows for more versatile test execution. As a result, we see an overfitting rate of 13.6% from USEagent on the SWE-verified dataset. In addition, the SWT and REPOTEST datasets show low numbers of overfitting, which further reduces the overall rate. Furthermore, we identified a set of 33 anomalies that are plausible, and do not overfit, yet they are different from the ground truth. One example anomaly includes the introduction of a new vector library for xarray (a vector library itself). Another example is an unorthodox class-inheritance check for scikit-learn - instead of utilizing pythons standard function isinstance, a predicate is applied to the objects obj.__classes__. These solutions are functionally correct but are unlikely to be accepted by developers, so we have marked them as anomalies.
Summary: RQ2 - Analysis of Plausible Solutions.  We manually inspected 218 plausible solutions from USEagent and report 10.5% overfitting, 1.3% memorization. Overfitting was most common in SWE-verified, and least common in REPOTEST and SWT.
@@ 6.3. RQ3: Self-Configuration
(a) Distribution of SWE Steps – the most common trajectory begins with reproduction, followed by code retrieval, editing, and review.
(b) Distribution of SWT Steps - the most common trajectory begins with test-case retrieval, followed by test-editing and execution.
@@
Table 4. Results of ablation study with USEagent on a statistically significant subset of USEbench, using claude-3-5-sonnet-20241022 as the backend model.
Dataset USEagent USEagent USEagent
original disable dynamic remove ExecuteTests
SWE-Ver 49.6% 46.2% (-7%) 49.6% (-0%)
SWT 40.6% 39.1% (-4%) 36.2% (-11%)
REPOCOD 8.7% 6.5% (-25%) 0% (-100%)
REPOTEST 25.0% 20.0% (-20%) 17.5% (-30%)
SWETRY 13.0% 8.7% (-33%) 8.7% (-33%)
Overall 34.9% 31.9% (-9%) 31.2% (-11%)
@@
Figure 4 shows the different choices of the Meta-Agent when facing different tasks (we present two for brevity). The histograms illustrate the actions taken at each step in the sequence (first, second, etc.). We observe some clear patterns in the different task types that match our intuition. For program repair tasks (i.e. in SWE-Verified), we see in Figure 4(a) that the first step prominently consists of generating a reproduction test, followed by retrieving relevant code context. Over the course of agent execution, the EditCode and ExecuteTests become more prominent, introducing and verifying code changes to the software. For regression testing (in SWT), Figure 4(b) shows that by-large the first step chosen is TestRetrieval, immediately followed by a (test-)code edit. The majority of USEagent trajectories on SWT tasks then consist of alternating test-changes and test-executions. Overall, the sequence of action invocation follows a certain pattern for each task type as shown in Figure 4, showing that USEagent can self-configure its workflow based on the task type. Across all benchmarks, USEagent first gathers information (using the retrieval actions or reproducers), converging towards alternating EditCode and ExecuteTests until either budgets are exhausted or Terminate is chosen. This behavior showcases the capability of emulating concepts like Test-Time Scaling (brown2024large) and command-discovery (bouzenia2024you) through the Meta-Agent.
Table 4 supports these findings through an ablation study. The disable dynamic column refers to USEagent with the dynamic decision making by MetaAgent disabled. Instead of letting the MetaAgent freely decide which action should be taken as the next step, we predefine the following workflows for the task types:
  • SWE/SWETRY: Reproduction -¿ CodeRetrieval -¿ EditCode -¿ (ReviewPatch -¿ EditCode)*
  • SWT: TestRetrieval -¿ CodeRetrieval -¿ EditCode -¿
    (ExecuteTests -¿ EditCode)*
  • REPOCOD: EditCode -¿ TestRetrieval -¿ EditCode -¿
    (ExecuteTests -¿ EditCode)*
  • REPOTEST: TestRetrieval -¿ EditCode -¿ (ExecuteTests -¿ EditCode)*
where * represents a loop, and the agent can decide when to break out the loop and end the workflow. We see that disabling dynamic decision making leads to reductions in efficacy from 4% to 33% across tasks. The MetaAgent in USEagent not only allows for it to handle multiple task types without hurting the efficacy compared to predefined workflows, but without it there would be an efficacy drop. This efficacy improvement (by the MetaAgent) is largely because the MetaAgent can help to recover from initial sub-optimal decisions (e.g. invoke retrieval actions towards the end of workflow to gather more context).
We performed another ablation study to examine the effect of an individual action. The column remove ExecuteTests in Table 4 shows the efficacy of USEagent after removing ExecuteTests, which is one of the actions newly introduced in this work. In task types that require repeated test executions to validate and refine solutions (e.g. REPOCOD and REPOTEST), removing ExecuteTests results in significant efficacy drops (-100% and -30%). On the other hand, there is no efficacy drop in SWE-bench tasks, since the MetaAgent could invoke alternative actions such as ReviewPatch to refine solutions. These results indicate that a single action can enhance effectiveness across multiple task types.
Summary: RQ3 - Self Configuration.USEagent is capable of choosing different action-patterns for different benchmarks. The self-configured workflow is as effective as predefined static workflows. In addition, a single action can help improve efficacy in multiple task types.
@@ 6.4. RQ4: Open Challenges in Agentic Systems
Based on our investigations, we identify several major challenges in the current unified software engineering agent. Firstly, when the task requires writing a large amount of code (e.g. feature development tasks), the agent often generates solutions that are “mostly correct”. This is reflected on our evaluation on the REPOCOD dataset, where the task is to implement a complex function based on natural language requirement, and the correctness is defined as passing a large number of hidden unit tests. We observe a number of solutions generated by the agent implement the required feature almost correctly, but fail a few hidden tests due to missing edge cases. The edge-case behaviors are sometimes not specified clearly in the natural-language task description. We attribute this challenge in part to the inherent ambiguity of natural language. A future agent may first resolve the ambiguity in natural language by transforming the task requirements into more “formal” artifacts such as specifications and test cases, and then generate solutions based on the formal specification. Another approach is to design human-agent interaction schemes to clarify ambiguity when needed.
@@
Secondly, the current agent lacks the capability of “backtracking” when the agent execution does not yield meaningful results on a given execution path. For example, when there is a partial patch at a particular program location, the agent is prone to continuously make small edits to the partial patch and execute tests to verify them. However, the agent may not make good progress with this approach, since it may be impossible to craft a good solution at the location of the partial patch. This issue is amplified in the SWETRY dataset, in which the task description comes with a partial patch. To tackle this issue, a future agent should employ a backtracking mechanism to discard unpromising partial solutions, and start over at a previous step. It could be possible to employ the recent reasoning LLMs to examine the agent execution trace and decide whether to backtrack, since reasoning models have shown backtracking behaviors in their thinking process (guo2025deepseek). Moreover, recent work (antoniades2024swesearch) on employing search algorithm over agent execution trajectories can help the agent escape from unpromising paths.
@@
Lastly, we still observe patch overfitting in tasks such as program repair and regression testing. Overfitting is a known problem for automated program repair (cacm19). Overfitting happens due to the generated patches passing given tests, but missing actual requirements. Thus, to avoid overfitting, if the tests are too specific, we need to generalize them. Agentic systems show promise for test generation (also reflected in this work), and the USEagent framework allows additional actions, such as test amplification or mutation testing, to be incorporated into the agent. Artifacts from these actions can be fed back into code generation, thereby reducing overfitting and enhancing trustworthiness.
@@
Summary: RQ4 - Open Challenges.  Challenges include exploring edge-case handling and alternative solutions in feature development, backtracking and discarding poor partial results in repair, and addressing patch overfitting.
@@ 7. Related Work
Table 5. Comparing USEagent with LLM agents for software engineering. All these agents take in natural language issues.
USEagent SWE-Agent OpenHands CodeAct AutoCode Rover Agentless MASAI Passerine CodeR
Is workflow predefined or dynamically constructed? dynamic dynamic dynamic predefined predefined predefined dynamic predefined
Handles only program repair or multiple task types? multiple multiple multiple program repair program repair program repair program repair program repair
What is the “unit of work” in the agent? SE action terminal-level command terminal-level command SE action SE action SE action SE action SE action
Multi-agent system? yes no no yes no yes no yes
@@
Recent Software Engineering Benchmarks fall into two broad categories: isolated coding exercises and repository-level tasks. Some coding datasets such as CodeXGlue (lu2021codexglue) come without an evaluation metric and form a training corpus; others such as HumanEval (chen2021codex), MBPP (austin2021program), ClassEval (du2023classeval) or CoderEval (yu2024codereval) provide a challenge accompanied by an evaluation harness (i.e., a test-suite). There are already ongoing efforts to provide a meta-benchmark for these tasks (yu2024humaneval; liu2024your), which commonly target LLMs directly rather than agentic systems. For USEbench, we select repository-level tasks to ensure challenges of uniform granularity. We opted against coding exercises, as many models have begun to saturate on such benchmarks (sallou2024breaking).
LLM Agents for Software Engineering
We consider a system agentic if the LLM exhibits a degree of autonomy and can interact with the environment through commands or other pre-defined tools. Within agentic systems, we distinguish terminal-based agents that interface LLMs directly with a console and minimal tooling (e.g., SWE-Agent (yang2024swe), Google Jules (google2024gemini2)), from approaches that integrate additional domain knowledge and techniques. These additions include specialized tools (such as spectrum-based fault localization in AutoCodeRover (zhang2024autocoderover)), self-reflective capabilities (such as the reviewer agent in SpecRover (ruan2024specrover)), and control logic (such as the finite-state machine in RepairAgent (bouzenia2024repairagent)). In this work, we have focused on AutoCodeRover (zhang2024autocoderover) and OpenHands (wang2024openhands), as both are widely used open-source projects. While the Aider project (AIDER2024) covers a well-suited scope involving multiple types of software engineering tasks, its workflow relies on human prompting. Passerine (rondon2025evaluating) by Google proposes a program repair framework inspired by SWE-Agent (yang2024swe) and incorporates a structure similar to our Meta-Agent, providing five pre-configured commands (and does not allow free execution of terminal commands, unlike SWE-Agent). Our approach differs in scope, as we target a broader range of tasks beyond program repair. MASAI (wadhwa2024masai) defines an action space for its central agent similar to our Meta-Agent, but is also limited to program repair. In contrast, we structure the essential artifacts involved in software debugging via a task state, incorporate a broader set of software engineering actions for orchestration, and apply the agentic framework to multiple task types. CodeR (chen2024coder) has a Meta-Agent component that pre-defines an execution plan, but does not adapt it dynamically as USEagent does. Table 5 compares USEagent with existing software engineering agents that take in natural language issues. USEagent distinguishes itself from the rest by dynamically orchestrating actions (each action is executed by sub-agents) to handle multiple types of software engineering tasks.
Figure 5. Average cost across Actions in USEagent.
@@ 8. Threats to validity
Within USEbench, we approximate testing capabilities using test coverage. While test coverage does not guarantee semantic coverage and can sometimes be achieved through various hacks, it remains a useful proxy in the absence of a better oracle. It captures several attributes we aim to evaluate: code changes must be correctly formatted and target the appropriate program locations. Moreover, upon inspection, many of the REPOTEST datapoints contain branches that handle edge-case behavior with errors, where syntactic coverage effectively reflects semantic testing.
Internal Validity: Data Leakage
It is possible that LLMs have been exposed to the projects under evaluation or even datapoints from the source benchmarks (sallou2024breaking). While fully preventing such exposure is beyond our capabilities and the scope of this work, we aim to provide an overview of potential memorization by manually inspecting a sample of plausible solutions.
@@ 9. Discussion
With Claude 3.5 Sonnet v2 as the backend model, USEagent costs were at an average of 3.65 USD, with notable differences across sub-benchmarks. Herein SWE scored lowest with 2.20 USD average, and REPOCOD highest at a mean of 7.66 USD. The high costs in REPOCOD originate from the iteration of patch generation and test execution. As shown in Figure 5, among all actions, EditCode is the main cost-driver, due to the large contexts provided through relevant code and tests. We calculate the Pearson correlation coefficient for the elapsed time and the costs, and find a strong correlation of $r=0.89$, implying correlation between computational (e.g. test execution) and model (e.g. reasoning) efforts. When using DeepSeek-V3 as backend model, the average cost of USEagent is 0.55 USD per datapoint, which is ~15% of the cost with Claude 3.5 Sonnet v2. Given the efficacy presented in Table 3, open-source models are cost-effective.
@@
We consider the approach of USEagent to be a promising avenue for solving any task that revolves around code as its primary artifact. Many tasks such as dependency management or project configuration can be incrementally tackled by introducing more actions into USEagent. For USEagent to become truly an AI Software Engineer, it must address tasks such as requirements engineering, data visualization, deployment, code review or even a-b-testing. The efforts presented in this work are the second step in agentic development towards an AI software engineer, the first being individual agentic systems for single tasks (like AutoCodeRover). With the USEagent we pave the way for a unified approach to interact autonomously with code regardless of the task. Going beyond the USEagent will involve studying the cooperative intelligence resulting from multiple USEagents and multiple human developers. This might establish the dynamics of future development teams.
Artifact Availability
We release the logs and evaluation of USEagent at https://doi.org/10.5281/zenodo.15023393, the source code of USEbench at https://github.com/nus-apr/USEbench, and the source code of USEagent at https://github.com/nus-apr/USEagent.
Acknowledgements.
This work was partially supported by a Singapore Ministry of Education (MoE) Tier 3 grant “Automated Program Repair”, MOE-MOET32021-0001, and the NUS Artificial Intelligence Institute (NAII) seed grant NAII-SF-2024-006. It was also supported in part by NSF 1901242 and 2006688 and a CFI fund.